40 men can do work in 25 days.how long it will take 25 men to do the work?
Answer by stanbon(75887)

40 men can do work in 25 days.how long it will take 25 men to do the work?
----
time is inversely related to # of men working
----
d = k/m
----
Solve for "k" using "40 men take 25 days"
25 = k/40
k = 1000
-----
Equation::
d = 1000/m
----
? how long it will take 25 men to do the work?
d = 1000/25 = 40 days
---------------
Cheers,
Stan H.
------------